### 30 Nov 2023 v1

# MLLMs-Augmented Visual-Language Representation Learning

Abstract
Visual-language pre-training (VLP) have achieved remarkable success in multi-modal tasks, largely attributed to the availability of large-scale image-text datasets. In this work, we demonstrate that multi-modal large language models (MLLMs) can enhance visual-language representation learning by improving data quality. Our approach is simple, utilizing MLLMs to extend multiple captions for each image. To prevent the bias that introduced by MLLMs’ hallucinations and intrinsic caption styles, we propose a “text shearing” to keep the lengths of extended captions identical to the originals. In image-text retrieval, our method consistently obtains 5.6 $\sim$ 35.0% and 16.8 $\sim$ 46.1% improvement on R@1 under the fine-tuning and zero-shot settings, respectively. Notably, our zero-shot results are comparable to fine-tuning on target datasets, which encourages more exploration on the versatile use of MLLMs.
1 Introduction
(a) Data Removing Ratios vs. Performance
(b) Data Rewriting Ratios vs. Performance
(c) The analysis of captioning the image using MiniGPT-4 and LLaVA.
Figure 1: The performance in (a) and (b) is zero-shot image-text retrieval results on MSCOCO using BLIP pretrained on CC3M. (c) shows that using a single model to generate captions can easily result in similar text structure and insufficient global considerations.
Visual-language pre-training (VLP) has exhibited remarkable achievements in image-text retrieval[25, 26], image classification[49, 66], vision question answering [3, 31], and image caption generation [32, 66]. The success of VLP can be attributed to the large-scale datasets that are collected from the Internet, such as CC3M [54], CC12M [8], YFCC100M [58], LAION400M [52], etc. However, most of these datasets include a non-negligible portion of noisy and mismatched image-text pairs [25, 60, 29], which largely affects the visual-language representation learning. One of the most straightforward approaches is: utilizing pre-trained models to recognize and remove mismatched pairs based on heuristic rules [17, 1, 7, 38].
These methods indeed reduce the influences of mismatched pairs. However, simply removing mismatched samples leads to a serious problem: the number of training pairs is also reduced. As illustrated in Fig. 0(a), the performance of image-text retrieval consistently drops a lot when the removing ratio is large. Most recently, several works [17, 69, 43] demonstrate that LLM and MLLM can be used as re-writers to improve caption quality without reducing the number of training pairs. Unfortunately, these re-writers inevitably introduce their caption styles, i.e. text structure, which might disrupt the distribution of the original captions and lead to difficulties in learning better visual-language representations.
To investigate the characteristics of previous rewriting works [17, 69, 43], we conduct experiments to evaluate its influences in Fig. 0(b). Experimental results empirically confirm that applying MLLMs to enhance the quality of visual-language datasets is a promising approach, i.e. improving the original performance largely. Nevertheless, excessive re-writing leads to non-trivial performance drops. Therefore, it is valuable to figure out what led to these drops and verify whether there exists bias from MLLMs’ captions.
We explore the underlying reasons for the bias by analyzing the generated captions’ text structures, attention focus, and word frequency. We utilize MiniGPT-4 and LLaVA to caption each image from CC3M, respectively. Here we use the same question {Describe the $\left\langle image\right\rangle$ in English:} to prompt the model and set the same maximum number of generated tokens. As shown in Fig. 0(c) and Fig. 3, we have the following observations. 1). MLLMs indeed have their inherent text structures. 2). Applying a single MLLM makes it hard to provide comprehensive captions. 3). There are great diversities between the word frequency statistics of different MLLMs and raw captions.
Considering the limitations of using a single model to caption the image in terms of diversity, we use multiple MLLMs to enrich visual-language associations from different perspectives, thereby enhancing visual-language representation learning. To mitigate the bias (from text structure and word frequency) introduced by MLLMs, the following specific designs are proposed. First, we reduce the length of captions generated by MLLMs to the same as the raw captions, which is called “text shearing”. It aims to mitigate the influences of intrinsic caption styles and hallucinations from MLLMs. Second, to obtain a comprehensive caption for each image, we keep the raw and extended captions from MLLMs simultaneously.
Our approach exhibits the following characteristics (or advantages): 1). It is compatible with multiple VLP methods like CLIP [49] and BLIP [31], demonstrating significant performance improvements across various downstream tasks. For example, in zero-shot image-text retrieval, our method obtains 16.8$\sim$ 46.1% improvement. 2). For image-text retrieval, without introducing additional training overhead, our zero-shot CLIP outperforms the model fine-tuned on the MSCOCO and Flickr30K respectively by 9.9% and 23.5%. 3). When scaling to large datasets like CC12M and YFCC15M, our method continues to deliver substantial performance improvements.
2 Related Work
2.1 Improving image-text datasets
Numerous studies [18, 42, 20] have emphasized the significance of high-quality image-text datasets in influencing the transfer performance of VLP models in downstream tasks. SemDeDup [1] enhances data efficiency and out-of-distribution performance by identifying and removing semantically duplicated data pairs at the embedding level. Cao et al. [7] introduced an approach to improve dataset quality by removing samples that contain text in images. T-MARS [38] masks the text region in the image and then filters out samples with low CLIP [49] similarity. These methods inevitably lose a lot of visual information when filtering out samples, so some methods try to obtain higher quality data by rewriting captions. Santurkar et al. [51] underscored the importance of synthetic captions and employed pre-trained language models to augment textual content. Gadre et al. [20] introduced the DataComp benchmark for multi-modal datasets. Fan et al. [17] leveraged the in-context learning capacity of large language models to rewrite captions, enriching the language structure while preserving core semantics. Zhu et al. [69] employed ChatGPT [45] and BLIP-2 [32] interactively to generate captions with rich visual information. Nguyen et al. [43] used BLIP-2 [32] to rewrite captions for low-matching image-text pairs and combined original and generated captions for training. Lai et al. [29] proposed enhancing visual information in captions by fusing the original caption with captions generated by LLaVA [36], leveraging large language models.
While approaches based on caption rewriting have yielded promising results, relying on a single model for rewriting introduces the model’s inherent bias and poses challenges in establishing a unified representation between vision and language. In contrast, our method introduces more accurate and diverse descriptions for a single image while retaining rich visual information.
2.2 Multimodal large language models
Existing Multimodal Large Language Models (MLLMs) primarily rely on three key technologies: Multimodal Instruction Tuning (M-IT), Multimodal In-Context Learning (M-ICL), and Multimodal Chain-of-Thought (M-CoT) [65]. M-IT facilitates strong transfer performance by fine-tuning the model on datasets with specific instruction formats. Notable models employing this technology include LLaMA-Adapter [67, 21], LLaVA [36, 35], MiniGPT-4 [70], InstructBLIP [13], Qwen-VL [5], and NExT-GPT [62]. M-ICL is a type of analogy learning from a limited number of samples. Models like Flamingo [3], Otter [30], and HuggingGPT [55] have been developed using this approach. M-CoT necessitates models to not only provide answers but also reasoning processes. Representative models include Multimodal-CoT [68] and Visual ChatGPT [61]. These three technologies are not mutually exclusive, and many models effectively combine multiple technologies. Since these MLLMs are often trained on billion-level dataset, they have extremely rich knowledge and excellent visual understanding and expression capabilities, which can be used to caption the images.
Figure 2: The illustration of our method. In terms of images, different MLLMs have different focuses. In terms of language, different MLLMs exhibit distinct idiomatic words and caption styles.
3 Method
3.1 Preliminaries
In this section, we briefly introduce the preliminaries of visual-language pre-training. CLIP [49] is a classic work that uses images and text for contrastive learning. It employs an encoder-only architecture, optimizing the encoder through a contrastive loss on image-text pairs. During training, when a batch of $N$ image-text pairs {$x_{I}$ , $x_{T}$ is sampled, the contrastive loss for an image can be defined as follows:
$L_{I}=\mathrm{clip\textunderscore contrast}(x_{I},x_{T})$ (1)
Where $\mathrm{clip\textunderscore contrast}$ is image-text contrastive loss introduced in  [49]. The loss for text is computed in the same manner and can be denoted as $L_{T}$. The total loss for training is $L=(L_{I}+L_{T})/2$.
BLIP [31] is an encoder-decoder based visual language pre-training architecture designed to enhance visual language understanding and generation. Its three main pre-training objectives (image-text contrastive loss (ITC), image-text matching loss (ITM), and language modeling loss (LM)) are jointly optimized during the training process. Specifically, ITC is similar to that in CLIP. ITM aims to capture the fine-grained alignment between vision and language. LM is used to optimize the decoder.
As highlighted in many works [66, 41, 33], visual-language pre-training heavily depends on extensive image-text datasets. The presence of low-quality image-text data can significantly compromise the model’s performance. Therefore, improving data quality has become an expected direction.
3.2 Overview
We demonstrate that different MLLMs can generate accurate and diverse captions in Fig. 2. The goal of this approach is to utilize the MLLMs to improve the data quality in the current image-text datasets. Our approach consists of two processes: multi-view caption extractor and text shearing. Specifically, for each image from an image-text dataset, we first introduce multiple advanced MLLMs to synthesize extended captions. Then, by comprehensively analyzing the characteristics of these extended captions, we have a key observation: MLLMs have their caption styles and might generate hallucinated words. Based on this observation, we propose text shearing to keep the length of extended captions the same as the raw one. After these processes, the raw and extended captions with corresponding images are used for standard visual-language pertaining methods.
3.3 Multi-View caption extractor
Given a dataset $\mathcal{T}=\{(\bm{x}_{I}^{i},x_{T}^{i})\}_{i=1}^{N}$, containing $N$ paired images $\bm{x}_{I}^{\cdot}$ and texts $x_{T}^{\cdot}$. We define a model pool $G=\{g_{1}\cdots g_{k}\cdots g_{K}\}$ that includes $K$ advanced MLLMs. For each image $\bm{x}_{I}^{i}$, $G$ is used to obtain rich captions. The operation can be formulated as follows,
$C^{i}=\{c_{1}^{i}\cdots c_{n}^{i}\cdots c_{N}^{i}\}=G(x_{I}^{i},Q)$ (2)
Where $C^{i}$ represents the extended captions for image $x_{I}^{i}$, and $Q$ denotes the question input to the MLLM. For different models in $G$, we use the same simple question template: {Describe the $\left\langle image\right\rangle$ in English:} to query the captions. The simple question has little impact on the diversity of answers, so we can obtain comprehensive captions of each image.
How to use these extended captions? After obtaining these extended captions, one of the most straightforward ideas is to replace the raw texts with a set of new captions for training the visual-language model. Its effectiveness has been evaluated in many previous works [69, 43]. However, we have a concern about the operation in terms of the difference between raw and new captions. To investigate this, we count the frequency of the most occurring nouns of the extended captions (extracted from images in CC3M by MiniGPT-4, Otter, Qwen-VL, and LLaVA-1.5, examples are shown in Fig. 2) and present the results in Fig. 3. One can find that different MLLMs have different common words in the generated captions.
Figure 3: Statistics of common nouns in captions generated by MLLMs.
Figure 4: Average length of captions generated using different MLLMs.
Based on the above analysis, we plan to use all the extended captions for training to improve the diversity of captions. However, there is another concern about the difference in the length of the raw and extended captions. We present the average length comparisons of these captions in Fig. 4. These captions are generated by MLLMs under default settings.It can be seen that the generated captions by MiniGPT-4 and LLaVA-1.5 are notably longer, i.e. 90 tokens. While raw caption is shorter, i.e. 15 tokens. To investigate the influence of varying caption lengths on pre-training, we visualize the longer caption generated by MiniGPT-4 in Figure 5. We find that longer captions indeed produce hallucinations that do not match the image, and these hallucinations often appear at the end of the caption, which has an impact on representation learning.
[…]
3.4 Text Shearing
In order to eliminate the impact of caption length, we propose text shearing to make the length of the extended captions as same as the raw ones. Specifically, we add a token number $T$ as a limit when using MLLMs to generate captions. In order to prevent the generated caption from having incomplete expressions, we intercept the first complete clause in the generated caption. Therefore, Eq. 2 can be rewritten as:
$C_{i}^{\prime}=\{c_{1}^{i^{\prime}}\cdots c_{n}^{i^{\prime}}\cdots c_{N}^{i^{\prime}}\}=G(x_{I}^{i},Q,T)$ (3)
For the setting of $T$, we use a simple but effective strategy. We set it to the average length of the original captions to mimic human annotation. By controlling the length of generated captions, we hope to effectively reduce the impact of MLLM’s hallucination and its intrinsic caption style.
Then, we jointly use original captions and newly generated captions to construct a new dataset. The new image-text pairs can be expressed as $\{x_{I},x_{T}^{\prime}\}=\{x_{I},c_{1}^{i^{\prime}}\},\{x_{I},c_{2}^{i^{\prime}}\}\cdots,\{x_{I},c_{k}^{i^{\prime}}\}$, where $k$ represents the number of MLLM. We use the enhanced dataset for subsequent standard visual-language pre-training. Taking CLIP as an example, the loss on the image can be written as:
$L_{I}=\mathrm{clip\textunderscore contrast}(x_{I},x_{T}^{\prime})$ (4)
The total loss $L=(L_{I}+L_{T})/2$. The parameters $\beta$ for CLIP is updated by minimizing $L$:
$\beta\leftarrow\operatorname*{arg\,min}_{\beta}L.$ (5)
Table 1: The results of zero-shot image-text retrieval and fine-tuning image-text retrieval on MSCOCO and Flickr30K.
Method Dataset MSCOCO (5K test set) Flickr30K (1K test set)
[…]
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
  CLIP [49] CC3M[54] 26.0 52.3 64.2 18.3 42.8 55.4 40.0 67.0 76.9 28.4 56.0 66.6
Ours 47.3 74.1 83.2 33.2 62.0 73.5 75.0 92.3 95.9 58.0 83.0 89.2
+21.3 +21.8 +19.0 +14.9 +19.2 +18.1 +35.0 +25.3 +19.0 +29.6 +27.0 +22.6
CC3M [54] 8.7 23.9 33.7 7.1 19.7 28.6 17.4 37.9 50.1 13.9 30.8 40.5
Ours 35.9 62.4 73.9 26.5 51.1 62.7 63.5 86.6 91.7 49.3 74.8 83.1
+27.2 +38.5 +40.2 +19.4 +31.4 +34.1 +46.1 +48.7 +41.6 +35.4 +44.0 +42.6
BLIP [31] CC3M[54] 67.5 88.9 93.7 50.9 76.9 85.2 88.5 98.5 99.5 74.0 92.4 95.7
Ours 76.2 93.0 96.4 57.6 82.3 89.3 94.1 99.4 99.8 82.2 95.7 97.9
+8.7 +4.1 +2.7 +6.7 +5.4 +4.1 +5.6 +0.9 +0.3 +8.2 +3.3 +2.2
CC3M [54] 36.3 62.6 73.6 28.6 52.9 64.2 62.1 87.4 91.9 51.2 75.5 82.7
Ours 59.3 81.5 88.9 45.2 71.0 80.2 85.5 97.1 98.8 72.0 90.3 94.3
+23.0 +18.9 +15.3 +16.8 +18.1 +16.0 +23.4 +9.7 +6.9 +20.8 +14.8 +11.6
Table 2: The zero-shot evaluation results of our method on 15 common classification datasets and ImageNet with CLIP pre-trained on CC3M and CC12M. The best results are highlighted in bold. On CC3M, our method demonstrates an average improvement of 13.4 on 15 datasets, and 13.1 on ImageNet. The different results on CLIP are due to the different training data downloaded from the Internet.
Data Model
Food-101
CIFAR-10
CIFAR-100
SUN397
Cars
Aircraft
DTD
Pets
Caltech-101
Flowers
STL-10
EuroSAT
RESISC45
GTSRB
Country211
Average
$\Delta$(%)
ImageNet
$\Delta$(%)
Model Architecture: ViT-B/16
CC3M CLIP 10.3 54.9 21.8 25.0 0.8 1.4 10.5 12.8 43.3 10.2 77.6 14.1 19.1 6.9 0.6 20.6 - 15.8 -
LaCLIP [17] 14.2 57.1 27.5 35.1 1.6 1.6 16.6 15.6 52.7 14.7 86.2 15.0 24.3 6.4 1.0 24.6 $\uparrow$4.0 21.5 $\uparrow$5.7
CLIP 9.2 28.4 9.8 21.9 1.1 1.1 8.3 7.6 40.7 8.9 70.4 13.2 14.8 5.6 0.4 16.1 - 11.9 -
Ours 18.7 58.4 32.4 43.8 3.9 1.5 20.2 32.1 63.5 17.5 87.3 25.1 23.1 13.0 2.0 29.5 $\uparrow$ 13.4 25.0 $\uparrow$ 13.1
CC12M CLIP 50.8 64.9 38.5 44.7 24.1 2.4 19.4 64.1 77.4 33.2 91.0 20.1 38.9 7.3 5.1 38.8 - 40.2 -
LaCLIP [17] 60.7 75.1 43.9 57.0 36.3 5.6 31.0 72.4 83.3 39.9 95.1 27.3 44.3 12.7 8.9 46.2 $\uparrow$7.4 48.4 $\uparrow$8.2
CLIP 45.9 65.7 33.4 44.7 18.4 2.9 18.5 54.8 72.6 30.4 89.5 23.2 28.5 9.3 4.8 36.2 - 37.3 -
Ours 60.9 83.0 55.4 59.4 24.1 3.2 30.7 64.8 79.3 36.0 95.3 40.5 45.6 25.4 5.8 47.3 $\uparrow$ 11.1 47.5 $\uparrow$ 10.2
4 Experiments
4.1 Pre-training Details
Our method is implemented in Pytorch [47] and trained on a node equipped with 8 NVIDIA A100 GPUs. For visual language pre-training, we follow the implementation of CLIP and BLIP [31] respectively. When pre-training CLIP, we use a randomly initialized ViT-B-16. For BLIP, we use a ViT-B-16 [16, 59] pre-trained on ImageNet [14] as the image encoder and $\text{Bert}_{\text{base}}$ [15] as the text encoder. Regarding image preprocessing, we employ random cropping to achieve $224\times 224$ during pre-training, and increase the image resolution to $384\times 384$ during fine-tuning. We apply RandomAugment [12] for image augmentation. For optimization, we utilize an AdamW [37] optimizer with a learning rate warm-up to 3e-4, combined with linear decay at a rate of 0.9. Our training batch size is set to 1280, and the number of training epochs is 4.
Regarding the usage of multi-modal large language models, we employed Mini-GPT4-Vicuna13B [70], Otter-Image-MPT7B  [30], Qwen-VL-Chat [5], and LLaVA-v1.5-13B [35]. During caption generation, we impose a maximum token limit of 30 to control the caption length. For the synthetic caption generation, we use beam search with number of beams equal to 1.
Table 3: The results of vision question answering, visual reasoning and image captioning of BLIP pre-trained on CC3M, CC12M and YFCC15M, respectively.
[…]
test-dev test-std val test dev test-P B@4 CIDEr CIDEr SPICE
CC3M [54] 71.5 71.8 28.9 24.0 76.0 76.2 36.6 119.9 84.4 12.5
Ours 75.6${}_{{\color[rgb]{0,0.8046875,0.421875}+4.1}}$ 75.6${}_{{\color[rgb]{0,0.8046875,0.421875}+3.8}}$ 32.0${}_{{\color[rgb]{0,0.8046875,0.421875}+3.1}}$ 31.4${}_{{\color[rgb]{0,0.8046875,0.421875}+7.4}}$ 80.1${}_{{\color[rgb]{0,0.8046875,0.421875}+4.1}}$ 79.3${}_{{\color[rgb]{0,0.8046875,0.421875}+3.1}}$ 37.6${}_{{\color[rgb]{0,0.8046875,0.421875}+1.0}}$ 125.6${}_{{\color[rgb]{0,0.8046875,0.421875}+5.7}}$ 97.0${}_{{\color[rgb]{0,0.8046875,0.421875}+12.6}}$ ${}_{{\color[rgb]{0,0.8046875,0.421875}+1.3}}$
CC12M [8] 73.5 73.5 31.5 31.0 78.7 79.0 37.2 123.1 99.4 13.5
Ours 77.0${}_{{\color[rgb]{0,0.8046875,0.421875}+3.5}}$ 77.1${}_{{\color[rgb]{0,0.8046875,0.421875}+3.6}}$ 32.0${}_{{\color[rgb]{0,0.8046875,0.421875}+0.5}}$ 31.6${}_{{\color[rgb]{0,0.8046875,0.421875}+0.6}}$ 81.2${}_{{\color[rgb]{0,0.8046875,0.421875}+2.5}}$ 81.9${}_{{\color[rgb]{0,0.8046875,0.421875}+2.9}}$ 39.1${}_{{\color[rgb]{0,0.8046875,0.421875}+1.9}}$ 130.1${}_{{\color[rgb]{0,0.8046875,0.421875}+7.0}}$ 103.6${}_{{\color[rgb]{0,0.8046875,0.421875}+4.2}}$ 14.3${}_{{\color[rgb]{0,0.8046875,0.421875}+0.8}}$
YFCC15M [58] 70.5 70.6 29.8 29.2 74.0 74.2 36.8 122.0 97.5 13.3
Ours 72.8${}_{{\color[rgb]{0,0.8046875,0.421875}+2.3}}$ 72.8+2.${}_{{\color[rgb]{0,0.8046875,0.421875}+2.2}}$ ${}_{{\color[rgb]{0,0.8046875,0.421875}+1.3}}$ 30.0${}_{{\color[rgb]{0,0.8046875,0.421875}+0.8}}$ 78.5${}_{{\color[rgb]{0,0.8046875,0.421875}+4.5}}$ 77.8${}_{{\color[rgb]{0,0.8046875,0.421875}+3.6}}$ 37.6${}_{{\color[rgb]{0,0.8046875,0.421875}+0.8}}$ 125.6${}_{{\color[rgb]{0,0.8046875,0.421875}+3.6}}$ 101.1${}_{{\color[rgb]{0,0.8046875,0.421875}+3.6}}$ 13.9${}_{{\color[rgb]{0,0.8046875,0.421875}+0.6}}$
4.2 Evaluation
We evaluate the performance on the image-text retrieval task in Table 1 and the image classification task in Table 2. In addition, we provide results on more tasks in Table 3.
Image-Text retrieval. Image-text retrieval serves as a crucial metric for evaluating the bridging capability between different modalities in a model. We pre-train CLIP and BLIP on CC3M, conducting image-to-text retrieval (TR) and text-to-image retrieval (IR) testing on the MSCOCO [34] and Flickr30K [48] datasets. Results in Table 1 demonstrate significant performance improvements in both zero-shot and fine-tuning retrievals for both CLIP and BLIP. Notably, our approach enables CLIP’s zero-shot retrieval performance to even exceed that achieved by fine-tuning on the target dataset. Specifically, when using CLIP for zero-sample retrieval on MSCOCO, the R@1 of TR and IR increase by 27.2 and 19.4, respectively. For zero-shot retrieval on Flickr30K, the R@1 of TR and IR also improves by 46.1 and 35.4, respectively. Similarly, our method also achieves an improvement of 16.8$\sim$23.4 when using BLIP for zero-shot retrieval. These significant performance enhancements indicate that the pretrained model obtains improved image-text representation.
Image Classification. We assess the performance of our method on image classification task. In Table 2, we pre-train a CLIP with a VIT-B-16 architecture on CC3M and CC12M respectively, followed by testing on sixteen common image classification datasets. Our method exhibits significant performance improvements across most datasets, as demonstrated. For the CLIP pre-trained on CC3M, there’s an average improvement of 13.4 on 15 datasets and a 13.1 improvement on ImageNet [14]. For the CLIP pre-trained on CC12M, there’s an average improvement of 11.1 and 10.2. These results highlight the effective enhancement of representation learning achieved by our method.
Vision Question Answering. Visual Question Answering (VQA) [4] is the task of providing answers based on an image and a question. We conduct tests on the VQAv2 [22], A-OKVQA [53] and OK-VQA [40] using BLIP pre-trained on CC3M, CC12M and YFCC15M, respectively. The results are presented in Table 3. Similar to BLIP [31], we also consider VQA as an answer generation task. The consistent performance improvements obtained through our method indicate that the model has acquired more robust visual-language representation from datasets enriched with MLLMs’ knowledge. Besides, the improvement on A-OKVQA and OK-VQA indicate that the model has more common sense and world knowledge. These results suggest an enhanced capability in visual understanding and language capabilities together.
[…]
CC3M [54] 26.0 46.3 58.0 7.0
[…]
CC12M [8] 34.0 56.7 68.5 4.0
[…]
YFCC15M [58] 28.2 48.0 63.2 6.0
[…]
Table 4: The results of performance on the video-language retrieval dataset MSRVTT.
Image Caption Generation. Image Captioning is the task of generating a text description of the image content given an image. We conduct tests on COCO and Nocaps [2] using the BLIP pre-trained on CC3M, CC12M and YFCC15M, respectively. Following the approach in [31] , we initially fine-tune the pre-trained model with the LM loss on COCO. To achieve better results, we also add ”a picture of” at the beginning of the prompt. The results are presented in Table 3, indicating improvements in BLEU@4 and CIDEr metrics. These findings suggest that our method enhances the model’s understanding of the relationship between text and images, resulting in higher similarity with human annotations during the captioning process.
(a) Token Length vs. Performance
(b) Batch size vs. Performance
(c) Model Numbers vs. Performance
(d) Epochs vs. Performance
Figure 6: The influence of the token length, batch size, model numbers and epochs on visual-language pre-training
Video-Language Task. Text-to-video retrieval is a metric for evaluating a model’s generalization ability in video-language tasks. We evaluate the effectiveness of our method on the MSRVTT dataset. Following the  [31], we fine-tune the model on COCO. For video input, we uniformly sample 8 frames from it to get a sequence. The results in the Table 3 demonstrate a stable performance improvement achieved by our method. This suggests that good visual-language representation learning may be key to video-text retrieval.
4.3 Analysis
Notably, as shown in Table 3, our method shows certain improvements when pre-training BLIP on CC12M [8] and YFCC15M [58]. This indicates that our method is scalable to larger datasets. It also emphasizes the significance of constructing well-represented image-text pairs for enhancing visual-language pre-training.
MLLM COCO (R@1) Flickr30k (R@1) Avg.
I2T T2I I2T T2I
Raw 36.3 28.6 62.1 51.2 44.6
MiniGPT-4 51.6 38.1 77.0 60.5 56.8
Otter 50.6 37.5 78.3 61.9 57.1
Qwen-VL 43.4 35.4 72.1 61.7 53.2
LLaVA-1.5 48.4 37.1 76.7 61.2 55.9
Table 5: Retrieval performance of models trained on captions rewritten using one MLLM.
4.4 Ablation Study
Figure 7: Long captions tend to have higher similarity due to the style limitations of the model itself. Similar content in captions is shown in red.
Caption Length. The length of the generated captions is one of the factors affecting visual-language pre-training. We utilize MiniGPT-4 to generate captions for CC3M with varying max token number limits. By training BLIP on captions of different lengths and conducting retrieval on MSCOCO , we present the results in Fig.5(a) and observe that as the length of the caption increases, the model’s performance tends to decrease.
This may be due to the fact that too long captions cause more similar text features between captions of different images. We illustrate this phenomenon in Fig. 7. When the caption length increases, the generated captions often become identical in many instances. This leads to different images easily mapping to the same text features, creating challenges in learning an effective representation.
Figure 8: Examples of image captioning using pre-trained BLIP without any fine-tuning. Newly introduced visual concepts are represented in red. It can be seen that BLIP pre-trained on our dataset has stronger visual concept capture capabilities.
Batch size. We explore the impact of batch size on training in the Fig. 5(b). Training models on synthetic captions requires relatively high batch sizes compared to using real captions annotated by humans. We attribute this to the fact that, with small batch size, the main gradient information obtained by the model is noisy, differing significantly from the true gradient. Conversely, with an increased batch size, the gradient of the sampled distribution becomes more similar to the gradient of the real distribution, enabling the model to generalize better.
Number of MLLMs. We explore the impact of the number of MLLMs on the transfer performance in the Fig. 5(c). We add four MLLMs{MiniGPT-4, Otter, Qwen-VL, and LLAVA-1.5}in order. Using pre-trained models for zero-shot retrieval on MSCOCO, we observe that, as the number of MLLMs increases, the performance gradually improves. However, the magnitude of this improvement is gradually decreasing, indicating that the one image’s caption information obtained from MLLM is reaching saturation.
Training epochs. We visualize the curve depicting the number of training epochs versus the model’s performance on the retrieval task in the Fig. 5(d). It is evident that our method achieves promising performance with only a small number of epochs. In the main experiment, we choose the epoch number of 4 to compare with the baseline without increasing the training cost. When further increasing the number of epochs, our method continues to show certain improvements.
Comparison using only one MLLM. We test the rewriting of captions using only one MLLM and utilize the generated captions for training. Captions are rewritten using MiniGPT-4, Otter, Qwen-VL, and LLaVA-1.5, and the results are presented in the Table 5. The results indicates that there is an upper limit to the performance improvement achievable by rewriting with only one model. Combining multiple MLLMs proves to be more effective in enhancing visual-language representation learning.
Figure 9: Cosine similarity distribution of images and text. Our method improves the average matching degree of image-text pairs.
Figure 10: Distribution of images and text. Our method aligns images and text better than the original distribution.
4.5 Visualization
Image and text distribution visualization. We visualize the feature distribution of image-text pairs from the original CC3M and ours dataset in the Fig. 10. It can be observed that the original CC3M dataset contains numerous discrete and unmatched image-text pairs. In contrast, our dataset exhibits a distribution where almost all images have corresponding texts. This distribution perspective explains how our method aligns more continuous images with discrete language, thereby enhancing visual-language representation learning.
Image captioning visualization. In the Fig. 8 we illustrate the difference in image captioning between models pre-trained using the original CC3M dataset and our improved dataset. without any fine-tuning, our model exhibits a significant improvement in the ability to recognize concepts in images. We provide more examples of image captioning in the appendix.
Cosine similarity visualization. We visualize the cosine similarity comparison between the original CC3M dataset and our dataset in the Fig. 9. We use a pre-trained CLIP to calculate the cosine distance between images and texts. It is evident that our dataset exhibits a higher average similarity than the original distribution.
5 Conclusion and Discussion
In this paper, we propose to augment visual-language representation learning by leveraging MLLMs. While retaining the rich visual information of the original dataset, we utilize multiple MLLMs to extend captions for each image. Additionally, we introduce “text shearing” to address issues of hallucinations and intrinsic caption style in synthetic captions. Validated across various VLP frameworks and datasets, our method significantly improves performance on numerous downstream tasks. This encourages further exploration in the utilization of MLLMs in future.
Limitations and future works Although our method has achieved remarkable results in enhancing visual-language representation learning, a certain proportion of noise still persists due to unreliable MLLM’s outputs. These noises limit further improvement of model’s performance. Future research could explore leveraging more powerful MLLMs and expanding to larger datasets.

### 13 Mar 2024 v3

@@
 Abstract
-Visual-language pre-training (VLP) have achieved remarkable success in multi-modal tasks, largely attributed to the availability of large-scale image-text datasets.
-In this work, we demonstrate that multi-modal large language models (MLLMs) can enhance visual-language representation learning by improving data quality.
-Our approach is simple, utilizing MLLMs to extend multiple captions for each image.
-To prevent the bias that introduced by MLLMs’ hallucinations and intrinsic caption styles, we propose a “text shearing” to keep the lengths of extended captions identical to the originals.
-In image-text retrieval, our method consistently obtains 5.6 $\sim$ 35.0% and 16.8 $\sim$ 46.1% improvement on R@1 under the fine-tuning and zero-shot settings, respectively.
-Notably, our zero-shot results are comparable to fine-tuning on target datasets, which encourages more exploration on the versatile use of MLLMs.
+Visual-language pre-training has achieved remarkable success in many multi-modal tasks, largely attributed to the availability of large-scale image-text datasets.
+In this work, we demonstrate that Multi-modal Large Language Models (MLLMs) can enhance visual-language representation learning by establishing richer image-text associations for image-text datasets.
+Our approach is simple, utilizing MLLMs to extend multiple diverse captions for each image.
+To prevent the bias introduced by MLLMs’ hallucinations and monotonous language styles, we propose “text shearing" to maintain the quality and availability of extended captions.
+In image-text retrieval, without introducing additional training cost, our method consistently obtains 5.6 $\sim$ 35.0% and 16.8 $\sim$ 46.1% improvement on Recall@1 under the fine-tuning and zero-shot settings, respectively.
+Notably, we obtain zero-shot results that are comparable to fine-tuning on target datasets, which encourages more exploration of the versatile use of MLLMs.
+The datasets and codes are available at: https://github.com/lyq312318224/MLLMs-Augmented.
+Keywords:
+Visual-language pre-training Representation learning
+**footnotetext: equal contributionfootnotetext: project leadfootnotetext: corresponding author
 1 Introduction
+(a) Data Removing Ratios vs. Performance
 (a) Data Removing Ratios vs. Performance
@@
 (c) The analysis of captioning the image using MiniGPT-4 and LLaVA.
-Figure 1: The performance in (a) and (b) is zero-shot image-text retrieval results on MSCOCO using BLIP pretrained on CC3M.
+Figure 1: The performance in (a) and (b) is zero-shot image-text retrieval results on MSCOCO using pretrained BLIP.
 (c) shows that using a single model to generate captions can easily result in similar text structure and insufficient global considerations.
-Visual-language pre-training (VLP) has exhibited remarkable achievements in image-text retrieval[25, 26], image classification[49, 66], vision question answering [3, 31], and image caption generation [32, 66].
-The success of VLP can be attributed to the large-scale datasets that are collected from the Internet, such as CC3M [54], CC12M [8], YFCC100M [58], LAION400M [52], etc.
+Visual-language pre-training has achieved remarkable success in image-text retrieval[25, 26], image classification[49, 66], vision question answering [3, 32], and image caption generation [31, 66].
+This success can be attributed to the large-scale datasets collected from the Internet, such as CC3M [54], CC12M [8], YFCC100M [58], LAION400M [52], etc.
 However, most of these datasets include a non-negligible portion of noisy and mismatched image-text pairs [25, 60, 29], which largely affects the visual-language representation learning.
 One of the most straightforward approaches is: utilizing pre-trained models to recognize and remove mismatched pairs based on heuristic rules [17, 1, 7, 38].
-These methods indeed reduce the influences of mismatched pairs.
-However, simply removing mismatched samples leads to a serious problem: the number of training pairs is also reduced.
-As illustrated in Fig.
-0(a), the performance of image-text retrieval consistently drops a lot when the removing ratio is large.
+These methods indeed reduce the influence of mismatched pairs.
+However, simply removing mismatched pairs leads to a serious problem: the number of training pairs is also reduced.
+As illustrated in Figure 0(b), the performance of image-text retrieval consistently drops a lot when the removal ratio is large.
 Most recently, several works [17, 70, 42] demonstrate that LLM and MLLM can be used as re-writers to improve caption quality without reducing the number of training pairs.
@@
 text structure, which might disrupt the distribution of the original captions and lead to difficulties in learning better visual-language representations.
-To investigate the characteristics of previous rewriting works [17, 69, 43], we conduct experiments to evaluate its influences in Fig.
-0(b).
+To investigate the characteristics of previous rewriting works [17, 70, 42], we conduct experiments to evaluate their influences in Figure 0(b).
 Experimental results empirically confirm that applying MLLMs to enhance the quality of visual-language datasets is a promising approach, i.e.
@@
 Nevertheless, excessive re-writing leads to non-trivial performance drops.
-Therefore, it is valuable to figure out what led to these drops and verify whether there exists bias from MLLMs’ captions.
+Therefore, it is valuable to figure out what led to these drops and verify whether there exists bias from synthetic captions.
 We explore the underlying reasons for the bias by analyzing the generated captions’ text structures, attention focus, and word frequency.
@@
 Here we use the same question {Describe the $\left\langle image\right\rangle$ in English:} to prompt the model and set the same maximum number of generated tokens.
-As shown in Fig.
-0(c) and Fig.
-3, we have the following observations.
+As shown in Figure 0(c) and Figure 3, we have the following observations.
 1).
-MLLMs indeed have their inherent text structures.
+MLLMs indeed have their inherent text structures and focus.
 2).
@@
 There are great diversities between the word frequency statistics of different MLLMs and raw captions.
-Considering the limitations of using a single model to caption the image in terms of diversity, we use multiple MLLMs to enrich visual-language associations from different perspectives, thereby enhancing visual-language representation learning.
-To mitigate the bias (from text structure and word frequency) introduced by MLLMs, the following specific designs are proposed.
-First, we reduce the length of captions generated by MLLMs to the same as the raw captions, which is called “text shearing”.
-It aims to mitigate the influences of intrinsic caption styles and hallucinations from MLLMs.
-Second, to obtain a comprehensive caption for each image, we keep the raw and extended captions from MLLMs simultaneously.
-Our approach exhibits the following characteristics (or advantages): 1).
-It is compatible with multiple VLP methods like CLIP [49] and BLIP [31], demonstrating significant performance improvements across various downstream tasks.
-For example, in zero-shot image-text retrieval, our method obtains 16.8$\sim$ 46.1% improvement.
+Considering the limitations of a single model in capturing diverse image captions, we employ multiple MLLMs to enrich visual-language associations from different perspectives, thereby improving visual-language representation learning during pre-training.
+To improve the quality and availability of synthetic captions, the following specific designs are proposed.
+First, we crop the captions generated by MLLMs to match the length of the original captions, which is called “text shearing”.
+This not only reduces the repetitive occurrence of common words in synthetic captions, alleviating the caption collapse problem identified in [60], but also preserves the semantic concepts closest to the image, thereby mitigating the impact of hallucinations typically appearing in the later part of the generated text, as verified in [69].
+Second, to obtain a set of comprehensive captions for each image, we keep the raw and extended captions from MLLMs simultaneously for standard visual-language pre-training.
+Our approach exhibits the following characteristics: 1).
+It is compatible with multiple visual-language pre-training frameworks like CLIP [49] and BLIP [32], demonstrating significant performance improvements across various downstream tasks without introducing additional training overhead.
+For example, in MSCOCO and Flickr30K’s zero-shot image-text retrieval, our method obtains 16.8$\sim$ 46.1% Recall@1 improvement.
+In zero-shot image classification, our method achieves an average performance improvement of 13.4 on 15 common classification datasets and 13.1 on ImageNet [14].
 2).
-For image-text retrieval, without introducing additional training overhead, our zero-shot CLIP outperforms the model fine-tuned on the MSCOCO and Flickr30K respectively by 9.9% and 23.5%.
-3).
+For image-text retrieval, our zero-shot CLIP outperforms the vanilla CLIP fine-tuned on the MSCOCO and Flickr30K respectively by 9.9% and 23.5%.
+(The above results are all based on pre-training on CC3M.) 3).
 When scaling to large datasets like CC12M and YFCC15M, our method continues to deliver substantial performance improvements.
@@
 2.1 Improving image-text datasets
-Numerous studies [18, 42, 20] have emphasized the significance of high-quality image-text datasets in influencing the transfer performance of VLP models in downstream tasks.
+Numerous studies [18, 43, 20] have emphasized the significance of high-quality image-text datasets in influencing the transfer performance of visual-language pre-training in downstream tasks.
 SemDeDup [1] enhances data efficiency and out-of-distribution performance by identifying and removing semantically duplicated data pairs at the embedding level.
@@
 T-MARS [38] masks the text region in the image and then filters out samples with low CLIP [49] similarity.
-These methods inevitably lose a lot of visual information when filtering out samples, so some methods try to obtain higher quality data by rewriting captions.
+These methods inevitably lose a lot of visual information when filtering out samples, so some methods try to obtain higher-quality data by rewriting captions.
 Santurkar et al.
@@
 M-ICL is a type of analogy learning from a limited number of samples.
-Models like Flamingo [3], Otter [30], and HuggingGPT [55] have been developed using this approach.
+Models like Flamingo [3], Otter [30], and HuggingGPT [55] are developed using this approach.
 M-CoT necessitates models to not only provide answers but also reasoning processes.
@@
 These three technologies are not mutually exclusive, and many models effectively combine multiple technologies.
-Since these MLLMs are often trained on billion-level dataset, they have extremely rich knowledge and excellent visual understanding and expression capabilities, which can be used to caption the images.
-Figure 2: The illustration of our method. In terms of images, different MLLMs have different focuses. In terms of language, different MLLMs exhibit distinct idiomatic words and caption styles.
+Since these MLLMs are often trained on the billion-level dataset, they have extremely rich knowledge and excellent visual understanding and expression capabilities, which can be used to caption the images.
+Figure 2: The illustration of our method. Different MLLMs jointly construct one-to-many image-text pairs with richer semantic associations.
 3 Method
@@
 In this section, we briefly introduce the preliminaries of visual-language pre-training.
-CLIP [49] is a classic work that uses images and text for contrastive learning.
+CLIP [49] is a classic work that uses image-text pairs for contrastive learning.
 It employs an encoder-only architecture, optimizing the encoder through a contrastive loss on image-text pairs.
@@
 The total loss for training is $L=(L_{I}+L_{T})/2$.
-BLIP [31] is an encoder-decoder based visual language pre-training architecture designed to enhance visual language understanding and generation.
+BLIP [32] is an encoder-decoder based visual-language pre-training architecture.
 Its three main pre-training objectives (image-text contrastive loss (ITC), image-text matching loss (ITM), and language modeling loss (LM)) are jointly optimized during the training process.
@@
 The presence of low-quality image-text data can significantly compromise the model’s performance.
-Therefore, improving data quality has become an expected direction.
+Therefore, improving the multi-modal dataset has become an expected direction.
 3.2 Overview
-We demonstrate that different MLLMs can generate accurate and diverse captions in Fig.
-2.
-The goal of this approach is to utilize the MLLMs to improve the data quality in the current image-text datasets.
+In this work, we employ multiple advanced MLLMs to augment the visual-language representation learning from a data-centric perspective.
+We demonstrate that different MLLMs can generate accurate and diverse captions in Figure 2.
+The goal of this approach is to utilize the MLLMs to establish richer image-text associations in the current datasets.
 Our approach consists of two processes: multi-view caption extractor and text shearing.
 Specifically, for each image from an image-text dataset, we first introduce multiple advanced MLLMs to synthesize extended captions.
-Then, by comprehensively analyzing the characteristics of these extended captions, we have a key observation: MLLMs have their caption styles and might generate hallucinated words.
-Based on this observation, we propose text shearing to keep the length of extended captions the same as the raw one.
-After these processes, the raw and extended captions with corresponding images are used for standard visual-language pertaining methods.
+Then, by comprehensively analyzing the characteristics of these extended captions, we have a key observation: MLLMs have their caption styles and might generate hallucinated content.
+Based on this observation, we propose text shearing to maintain the quality and availability of extended captions.
+After these processes, the raw and extended captions with corresponding images are jointly used for standard visual-language pre-training.
 3.3 Multi-View caption extractor
@@
 How to use these extended captions?
-After obtaining these extended captions, one of the most straightforward ideas is to replace the raw texts with a set of new captions for training the visual-language model.
+After obtaining these extended captions, one of the most straightforward ideas is to replace a set of raw captions with new captions for training.
 Its effectiveness has been evaluated in many previous works [70, 42].
 However, we have a concern about the operation in terms of the difference between raw and new captions.
-To investigate this, we count the frequency of the most occurring nouns of the extended captions (extracted from images in CC3M by MiniGPT-4, Otter, Qwen-VL, and LLaVA-1.5, examples are shown in Fig.
-2) and present the results in Fig.
-3.
-One can find that different MLLMs have different common words in the generated captions.
+To investigate this, we count the frequency of the most occurring nouns of the extended captions (extracted from images in CC3M by MiniGPT-4, Otter, Qwen-VL, and LLaVA-1.5, examples are shown in Figure 2) and present the results in Figure 3.
+One can easily find that different MLLMs output different common words in the generated captions.
+These diverse words can greatly expand the linguistic concepts of the dataset.
 Figure 3: Statistics of common nouns in captions generated by MLLMs.
-Figure 4: Average length of captions generated using different MLLMs.
+(a) Average length of captions generated using different MLLMs.
+[…]
+Figure 4: Analysis of extended captions.
 Based on the above analysis, we plan to use all the extended captions for training to improve the diversity of captions.
 However, there is another concern about the difference in the length of the raw and extended captions.
-We present the average length comparisons of these captions in Fig.
-4.
+We present the average length comparisons of these captions in Figure 3(a).
 These captions are generated by MLLMs under default settings.It can be seen that the generated captions by MiniGPT-4 and LLaVA-1.5 are notably longer, i.e.
@@
 15 tokens.
-To investigate the influence of varying caption lengths on pre-training, we visualize the longer caption generated by MiniGPT-4 in Figure 5.
-We find that longer captions indeed produce hallucinations that do not match the image, and these hallucinations often appear at the end of the caption, which has an impact on representation learning.
-[…]
+To investigate the influence of varying caption lengths on pre-training, we visualize the longer caption generated by MiniGPT-4 in Figure 3(b).
+We find that longer captions indeed produce hallucinations that do not match the image, and these hallucinations often appear in the later part of the generated text, which has an impact on representation learning.
+The same conclusion is also verified in  [69].
+It states that as generation progresses, the accumulation of past hallucinatory information and uncertainty can derail the model and lead to new hallucinations.
 3.4 Text Shearing
-In order to eliminate the impact of caption length, we propose text shearing to make the length of the extended captions as same as the raw ones.
+To eliminate the impact of excessive caption length, we propose text shearing to make the length of the extended captions as same as the raw ones.
 Specifically, we add a token number $T$ as a limit when using MLLMs to generate captions.
-In order to prevent the generated caption from having incomplete expressions, we intercept the first complete clause in the generated caption.
+To prevent the generated captions from having incomplete expressions, we intercept the first complete clause in the generated caption.
 Therefore, Eq.
@@
 We set it to the average length of the original captions to mimic human annotation.
-By controlling the length of generated captions, we hope to effectively reduce the impact of MLLM’s hallucination and its intrinsic caption style.
+By controlling the length of generated captions, we hope to effectively reduce the impact of MLLMs’ hallucinations and their monotonous language styles.
 Then, we jointly use original captions and newly generated captions to construct a new dataset.
-The new image-text pairs can be expressed as $\{x_{I},x_{T}^{\prime}\}=\{x_{I},c_{1}^{i^{\prime}}\},\{x_{I},c_{2}^{i^{\prime}}\}\cdots,\{x_{I},c_{k}^{i^{\prime}}\}$, where $k$ represents the number of MLLM.
+The new image-text pairs can be expressed as $\{x_{I},x_{T}^{\prime}\}=\{x_{I},c_{1}^{i^{\prime}}\},\{x_{I},c_{2}^{i^{\prime}}\}\cdots,\{x_{I},c_{k}^{i^{\prime}}\}$, where $k$ represents the number of MLLMs.
 We use the enhanced dataset for subsequent standard visual-language pre-training.
@@
 $\beta\leftarrow\operatorname*{arg\,min}_{\beta}L.$ (5)
+4 Experiments
+4.1 Pre-training Details
+Our method is implemented in Pytorch [47] and trained on a node equipped with 8 NVIDIA A100 GPUs.
+For visual language pre-training, we follow the implementation of CLIP [49] and BLIP [32] respectively.
+For CLIP, we train the model based on the open source code OpenCLIP [24].
+During training, our batch size is set to 320.
+The number of epochs is set to 6 to better align the training overhead.
+The model architecture is set as VIT-B-16.
+Other parameters are the same as the default values in OpenCLIP [24].
+For BLIP, we use a ViT-B-16 [16, 59] pre-trained on ImageNet [14] as the image encoder and $\text{Bert}_{\text{base}}$ [15] as the text encoder.
+Regarding image preprocessing, we employ random cropping to achieve $224\times 224$ during pre-training and increase the image resolution to $384\times 384$ during fine-tuning.
+We apply RandomAugment [12] for image augmentation.
+For optimization, we utilize an AdamW [37] optimizer with a learning rate warm-up to 3e-4, combined with linear decay at a rate of 0.9.
+Our training batch size is set to 1280, and the number of training epochs is 4.
+Regarding the usage of multi-modal large language models, we employed Mini-GPT4-Vicuna13B [71], Otter-Image-MPT7B  [30], Qwen-VL-Chat [5], and LLaVA-v1.5-13B [35].
+During caption generation, we impose a maximum token limit of 30 to control the caption length.
+For the synthetic caption generation, we use beam search with the number of beams equal to 1.
+We use CC3M [54], CC12M [8] and YFCC15M as pre-training datasets.
+Among them, YFCC15M is a subset of YFCC100M [58] that contains English descriptions of images.
+Due to the inaccessibility of some images, the version of the dataset we used contains fewer images than the original dataset.
+Among them, on CC3M, compared with the 3.3M version of the original dataset, the version we used contains 2.6M image-text pairs; on CC12M, compared with the 12.4M version of the original dataset, the version we used contains 11.1M image-text pairs; On YFCC15M, compared to the 15M of the original dataset, our version contains 14.8M image-text pairs.
 Table 1: The results of zero-shot image-text retrieval and fine-tuning image-text retrieval on MSCOCO and Flickr30K.
@@
 +23.0 +18.9 +15.3 +16.8 +18.1 +16.0 +23.4 +9.7 +6.9 +20.8 +14.8 +11.6
+4.2 Evaluation
+We evaluate the performance of the multi-modal task in Table 1 and Table 4. We also provide results on the image classification task in Table 2 and Table 3.
+Image-Text Retrieval.
+Image-text retrieval serves as a crucial metric for evaluating the bridging capability between different modalities in a model.
+We pre-train CLIP and BLIP on CC3M, conducting image-to-text retrieval (TR) and text-to-image retrieval (IR) testing on the MSCOCO [34] and Flickr30K [48] datasets.
+Results in Table 1 demonstrate significant performance improvements in both zero-shot and fine-tuning retrievals for both CLIP and BLIP.
+Notably, our approach enables CLIP’s zero-shot retrieval performance to outperform models finetuned on the target dataset.
+Specifically, when using CLIP for zero-shot retrieval on MSCOCO, the R@1 of TR and IR increase by 27.2 and 19.4, respectively.
+For zero-shot retrieval on Flickr30K, the R@1 of TR and IR also improves by 46.1 and 35.4, respectively.
+Similarly, our method also achieves an improvement of 16.8$\sim$23.4 when using BLIP for zero-shot retrieval.
+These significant performance improvements indicate that the pre-trained model learns better image-text representation.
 Table 2: The zero-shot evaluation results of our method on 15 common classification datasets and ImageNet with CLIP pre-trained on CC3M and CC12M.
 The best results are highlighted in bold.
-On CC3M, our method demonstrates an average improvement of 13.4 on 15 datasets, and 13.1 on ImageNet.
-The different results on CLIP are due to the different training data downloaded from the Internet.
+On CC3M, our method demonstrates an average improvement of 13.4 on 15 datasets and 13.1 on ImageNet.
+The different results on vanilla CLIP are due to the different training data downloaded from the Internet.
 Data Model
@@
 Ours 60.9 83.0 55.4 59.4 24.1 3.2 30.7 64.8 79.3 36.0 95.3 40.5 45.6 25.4 5.8 47.3 $\uparrow$ 11.1 47.5 $\uparrow$ 10.2
-4 Experiments
-4.1 Pre-training Details
-Our method is implemented in Pytorch [47] and trained on a node equipped with 8 NVIDIA A100 GPUs.
-For visual language pre-training, we follow the implementation of CLIP and BLIP [31] respectively.
-When pre-training CLIP, we use a randomly initialized ViT-B-16.
-For BLIP, we use a ViT-B-16 [16, 59] pre-trained on ImageNet [14] as the image encoder and $\text{Bert}_{\text{base}}$ [15] as the text encoder.
-Regarding image preprocessing, we employ random cropping to achieve $224\times 224$ during pre-training, and increase the image resolution to $384\times 384$ during fine-tuning.
-We apply RandomAugment [12] for image augmentation.
-For optimization, we utilize an AdamW [37] optimizer with a learning rate warm-up to 3e-4, combined with linear decay at a rate of 0.9.
-Our training batch size is set to 1280, and the number of training epochs is 4.
-Regarding the usage of multi-modal large language models, we employed Mini-GPT4-Vicuna13B [70], Otter-Image-MPT7B  [30], Qwen-VL-Chat [5], and LLaVA-v1.5-13B [35].
-During caption generation, we impose a maximum token limit of 30 to control the caption length.
-For the synthetic caption generation, we use beam search with number of beams equal to 1.
-Table 3: The results of vision question answering, visual reasoning and image captioning of BLIP pre-trained on CC3M, CC12M and YFCC15M, respectively.
+Image Classification.
+We evaluate the performance of our method on the task of image classification.
+In Table 2, we pre-train a CLIP with a VIT-B-16 architecture on CC3M and CC12M respectively, followed by testing on sixteen common image classification datasets.
+We compare our results with  [17].
+Our method exhibits significant performance improvements across most datasets, as demonstrated.
+For the CLIP pre-trained on CC3M, there’s an average improvement of 13.4 on 15 datasets and a 13.1 improvement on ImageNet [14].
+For the CLIP pre-trained on CC12M, there’s an average improvement of 11.1 and 10.2 respectively.
+These results highlight the effective enhancement of representation learning achieved by our method.
+Table 3: The linear probing results of our method on 15 common classification datasets with CLIP pre-trained on CC3M and CC12M.
+On CC3M, our method demonstrates an average improvement of 7.9 on 15 datasets.
+Data Model
+Food-101
+CIFAR-10
+CIFAR-100
+SUN397
+Cars
+Aircraft
+DTD
+Pets
+Caltech-101
+Flowers
+STL-10
+EuroSAT
+RESISC45
+GTSRB
+Country211
+Average
+$\Delta$(%)
+Model Architecture: ViT-B/16
+CC3M CLIP 62.6 86.8 68.1 58.5 32.8 40.9 63.4 69.6 82.0 89.4 91.7 95.9 89.0 71.9 13.3 67.7 -
+LaCLIP [17] 63.8 87.7 69.5 60.2 32.4 42.7 64.0 71.1 83.3 90.2 93.4 95.8 89.7 74.6 13.2 68.8 $\uparrow$1.1
+CLIP 52.8 79.7 58.8 52.4 21.9 26.4 53.1 56.6 77.5 76.6 85.7 93.9 84.5 66.7 8.7 59.7 -
+Ours 64.0 87.7 68.5 59.1 34.5 32.1 60.4 73.3 85.5 83.6 92.6 95.3 87.9 78.4 10.6 67.6 $\uparrow$ 7.9
+CC12M CLIP 81.6 93.8 79.3 72.0 75.1 52.6 75.6 86.2 92.2 95.3 97.3 96.7 93.1 80.6 19.7 79.4 -
+LaCLIP [17] 82.9 94.7 79.7 73.8 79.9 54.5 75.7 87.7 93.0 96.4 98.0 96.4 93.0 81.9 19.7 80.5 $\uparrow$1.1
+CLIP 73.9 89.9 71.2 68.9 62.8 35.6 67.1 81.1 90.3 87.7 95.0 95.6 89.1 77.4 13.8 73.3 -
+Ours 80.6 93.9 78.1 72.1 62.1 35.6 72.1 82.7 90.6 87.9 97.3 95.7 91.1 84.8 13.8 75.9 $\uparrow$ 2.6
+Linear-Probing.
+We also explore the linear-probing performance in image classification.
+We use CLIP pre-trained on CC3M and CC12M for evaluation respectively.
+Similar to the zero-shot image classification, we compare our method with [17] across 15 common datasets.
+The results in Table 3 show that our method also has certain improvements for linear probing.
+This shows that rich text concepts contribute to the effective training of visual encoders.
+As the visual encoder learns each image, it benefits from supervision through multiple accurate and diverse captions.
+This allows it to efficiently acquire a universal representation of the image, leading to a more effective characterization of the image and text embedding space.
+Vision Question Answering.
+Visual Question Answering (VQA) [4] is the task of providing answers based on an image and a question.
+We conduct tests on the VQAv2 [22], A-OKVQA [53] and OK-VQA [40] using BLIP pre-trained on CC3M, CC12M and YFCC15M, respectively.
+The results are presented in Table 4.
+Similar to BLIP [32], we also consider VQA as an answer-generation task.
+The consistent performance improvements obtained through our method indicate that the model has acquired a more robust visual-language representation from datasets enriched with MLLMs’ knowledge.
+Besides, the improvement on A-OKVQA and OK-VQA indicates that the model has more common sense and world knowledge.
+These results suggest an enhanced capability in visual understanding and language capabilities together.
+Table 4: The results of vision question answering, visual reasoning and image captioning of BLIP pre-trained on CC3M, CC12M, and YFCC15M, respectively.
 […]
@@
 Ours 72.8${}_{{\color[rgb]{0,0.8046875,0.421875}+2.3}}$ 72.8+2.${}_{{\color[rgb]{0,0.8046875,0.421875}+2.2}}$ ${}_{{\color[rgb]{0,0.8046875,0.421875}+1.3}}$ 30.0${}_{{\color[rgb]{0,0.8046875,0.421875}+0.8}}$ 78.5${}_{{\color[rgb]{0,0.8046875,0.421875}+4.5}}$ 77.8${}_{{\color[rgb]{0,0.8046875,0.421875}+3.6}}$ 37.6${}_{{\color[rgb]{0,0.8046875,0.421875}+0.8}}$ 125.6${}_{{\color[rgb]{0,0.8046875,0.421875}+3.6}}$ 101.1${}_{{\color[rgb]{0,0.8046875,0.421875}+3.6}}$ 13.9${}_{{\color[rgb]{0,0.8046875,0.421875}+0.6}}$
-4.2 Evaluation
-We evaluate the performance on the image-text retrieval task in Table 1 and the image classification task in Table 2. In addition, we provide results on more tasks in Table 3.
-Image-Text retrieval.
-Image-text retrieval serves as a crucial metric for evaluating the bridging capability between different modalities in a model.
-We pre-train CLIP and BLIP on CC3M, conducting image-to-text retrieval (TR) and text-to-image retrieval (IR) testing on the MSCOCO [34] and Flickr30K [48] datasets.
-Results in Table 1 demonstrate significant performance improvements in both zero-shot and fine-tuning retrievals for both CLIP and BLIP.
-Notably, our approach enables CLIP’s zero-shot retrieval performance to even exceed that achieved by fine-tuning on the target dataset.
-Specifically, when using CLIP for zero-sample retrieval on MSCOCO, the R@1 of TR and IR increase by 27.2 and 19.4, respectively.
-For zero-shot retrieval on Flickr30K, the R@1 of TR and IR also improves by 46.1 and 35.4, respectively.
-Similarly, our method also achieves an improvement of 16.8$\sim$23.4 when using BLIP for zero-shot retrieval.
-These significant performance enhancements indicate that the pretrained model obtains improved image-text representation.
-Image Classification.
-We assess the performance of our method on image classification task.
-In Table 2, we pre-train a CLIP with a VIT-B-16 architecture on CC3M and CC12M respectively, followed by testing on sixteen common image classification datasets.
-Our method exhibits significant performance improvements across most datasets, as demonstrated.
-For the CLIP pre-trained on CC3M, there’s an average improvement of 13.4 on 15 datasets and a 13.1 improvement on ImageNet [14].
-For the CLIP pre-trained on CC12M, there’s an average improvement of 11.1 and 10.2.
-These results highlight the effective enhancement of representation learning achieved by our method.
-Vision Question Answering.
-Visual Question Answering (VQA) [4] is the task of providing answers based on an image and a question.
-We conduct tests on the VQAv2 [22], A-OKVQA [53] and OK-VQA [40] using BLIP pre-trained on CC3M, CC12M and YFCC15M, respectively.
-The results are presented in Table 3.
-Similar to BLIP [31], we also consider VQA as an answer generation task.
-The consistent performance improvements obtained through our method indicate that the model has acquired more robust visual-language representation from datasets enriched with MLLMs’ knowledge.
-Besides, the improvement on A-OKVQA and OK-VQA indicate that the model has more common sense and world knowledge.
-These results suggest an enhanced capability in visual understanding and language capabilities together.
 […]
+Image Captioning.
+Image Captioning is the task of generating a text description of the image content given an image.
+We conduct tests on COCO and Nocaps [2] using the BLIP pre-trained on CC3M, CC12M, and YFCC15M, respectively.
+Following the approach in [32], we initially fine-tune the pre-trained model with the LM loss on COCO.
+To achieve better results, we also add "a picture of" at the beginning of the prompt.
+The results are presented in Table 4, indicating improvements in BLEU@4 and CIDEr metrics.
+These findings suggest that our method enhances the model’s understanding of the relationship between texts and images, resulting in higher similarity with human annotations during the captioning process.
 […]
@@
 […]
-Table 4: The results of performance on the video-language retrieval dataset MSRVTT.
-Image Caption Generation.
-Image Captioning is the task of generating a text description of the image content given an image.
-We conduct tests on COCO and Nocaps [2] using the BLIP pre-trained on CC3M, CC12M and YFCC15M, respectively.
-Following the approach in [31] , we initially fine-tune the pre-trained model with the LM loss on COCO.
-To achieve better results, we also add ”a picture of” at the beginning of the prompt.
-The results are presented in Table 3, indicating improvements in BLEU@4 and CIDEr metrics.
-These findings suggest that our method enhances the model’s understanding of the relationship between text and images, resulting in higher similarity with human annotations during the captioning process.
+Table 5: The retrieval performance on the video-language retrieval dataset MSRVTT. R@K stands for recall at K and MdR stands for median rank.
+Video-Language Task.
+Text-to-video retrieval is a metric for evaluating a model’s generalization ability in video-language tasks.
+We evaluate our method on the MSRVTT dataset.
+Following the  [32], we fine-tune the model on COCO.
+For video input, we uniformly sample 8 frames from it to get a sequence.
+The results in Table 4 demonstrate a stable performance improvement achieved by our method.
+This suggests that robust visual-language representation learning may be key to video-text retrieval.
+4.3 Analysis
+Scaling Ability.
+Notably, as shown in Table 4, our method shows certain improvements when pre-training BLIP on CC12M [8] and YFCC15M [58].
+This indicates that our method is scalable to larger datasets to some extent.
+It also emphasizes the significance of constructing well-represented image-text pairs for enhancing visual-language pre-training.
+Training Cost.
+To ensure a fair comparison with baseline methods, we carefully craft the training schedule.
+Considering that the number of image-text pairs in the augmented dataset is $k$ times the number of image-text pairs in the original dataset, we adjust the pre-training epochs to be $1/k$ of the original epochs.
+By doing this, we avoid introducing additional training overhead.
 (a) Token Length vs. Performance
@@
 (d) Epochs vs. Performance
-Figure 6: The influence of the token length, batch size, model numbers and epochs on visual-language pre-training
-Video-Language Task.
-Text-to-video retrieval is a metric for evaluating a model’s generalization ability in video-language tasks.
-We evaluate the effectiveness of our method on the MSRVTT dataset.
-Following the  [31], we fine-tune the model on COCO.
-For video input, we uniformly sample 8 frames from it to get a sequence.
-The results in the Table 3 demonstrate a stable performance improvement achieved by our method.
-This suggests that good visual-language representation learning may be key to video-text retrieval.
-4.3 Analysis
-Notably, as shown in Table 3, our method shows certain improvements when pre-training BLIP on CC12M [8] and YFCC15M [58].
-This indicates that our method is scalable to larger datasets.
-It also emphasizes the significance of constructing well-represented image-text pairs for enhancing visual-language pre-training.
+Figure 5: The influence of the token length, batch size, model numbers, and epochs on the performance of visual-language pre-training.
+4.4 Ablation Study
+Caption Length.
+The length of the generated captions is one of the factors affecting visual-language pre-training.
+We utilize MiniGPT-4 to generate captions for CC3M with varying max token number limits.
+By training BLIP on captions of different lengths and conducting retrieval on MSCOCO, we present the results in Figure4(a) and observe that as the length of the caption increases, the model’s performance tends to decrease.
+This may be because too-long captions
+Figure 6: Long captions for different images tend to be similar due to the intrinsic inference manner of the model. Similar content in captions is shown in red.
+cause more similar text features between captions of different images.
+We illustrate this phenomenon in Figure 6.
+When the caption length increases, the generated captions often become identical in many instances.
+This leads to different images easily mapping to the same text features, creating challenges in learning an effective representation.
+Figure 7: Examples of image captioning using pre-trained BLIP without any fine-tuning.
+Newly introduced visual concepts are represented in red.
+It can be seen that BLIP pre-trained on our dataset has stronger visual concept capture capabilities.
+Batch Size.
+We explore the impact of batch size on training in Figure 4(b).
+Training models on synthetic captions requires a relatively high batch size compared to using captions annotated by humans.
+We attribute this to the fact that, with a small batch size, the main gradient information obtained is noisy, differing significantly from the true gradient.
+Conversely, with an increased batch size, the gradient of the sampled distribution becomes more similar to the gradient of the real distribution, enabling the model to generalize better.
+Number of MLLMs.
+We explore the impact of the number of MLLMs on the transfer performance in Figure 4(c).
+We add four MLLMs{MiniGPT-4, Otter, Qwen-VL, and LLAVA-1.5}in order.
+Using pre-trained models for zero-shot retrieval on MSCOCO, we observe that, as the number of MLLMs increases, the performance gradually improves.
+However, the magnitude of this improvement is gradually decreasing, indicating that the one image’s caption information obtained from MLLM is reaching saturation.
+Training Epochs.
+We visualize the curve depicting the number of training epochs versus the model’s performance on the retrieval task in Figure 4(d).
+It is evident that our method achieves promising performance with only a small number of epochs.
+In the main experiment, we choose epoch number 4 to compare with the baseline without increasing the training cost.
+When further increasing the number of epochs, our method continues to show certain improvements.
 MLLM COCO (R@1) Flickr30k (R@1) Avg.
@@
 LLaVA-1.5 48.4 37.1 76.7 61.2 55.9
-Table 5: Retrieval performance of models trained on captions rewritten using one MLLM.
-4.4 Ablation Study
-Figure 7: Long captions tend to have higher similarity due to the style limitations of the model itself. Similar content in captions is shown in red.
-Caption Length.
-The length of the generated captions is one of the factors affecting visual-language pre-training.
-We utilize MiniGPT-4 to generate captions for CC3M with varying max token number limits.
-By training BLIP on captions of different lengths and conducting retrieval on MSCOCO , we present the results in Fig.5(a) and observe that as the length of the caption increases, the model’s performance tends to decrease.
-This may be due to the fact that too long captions cause more similar text features between captions of different images.
-We illustrate this phenomenon in Fig.
-7.
-When the caption length increases, the generated captions often become identical in many instances.
-This leads to different images easily mapping to the same text features, creating challenges in learning an effective representation.
-Figure 8: Examples of image captioning using pre-trained BLIP without any fine-tuning.
-Newly introduced visual concepts are represented in red.
-It can be seen that BLIP pre-trained on our dataset has stronger visual concept capture capabilities.
-Batch size.
-We explore the impact of batch size on training in the Fig.
-5(b).
-Training models on synthetic captions requires relatively high batch sizes compared to using real captions annotated by humans.
-We attribute this to the fact that, with small batch size, the main gradient information obtained by the model is noisy, differing significantly from the true gradient.
-Conversely, with an increased batch size, the gradient of the sampled distribution becomes more similar to the gradient of the real distribution, enabling the model to generalize better.
-Number of MLLMs.
-We explore the impact of the number of MLLMs on the transfer performance in the Fig.
-5(c).
-We add four MLLMs{MiniGPT-4, Otter, Qwen-VL, and LLAVA-1.5}in order.
-Using pre-trained models for zero-shot retrieval on MSCOCO, we observe that, as the number of MLLMs increases, the performance gradually improves.
-However, the magnitude of this improvement is gradually decreasing, indicating that the one image’s caption information obtained from MLLM is reaching saturation.
-Training epochs.
-We visualize the curve depicting the number of training epochs versus the model’s performance on the retrieval task in the Fig.
-5(d).
-It is evident that our method achieves promising performance with only a small number of epochs.
-In the main experiment, we choose the epoch number of 4 to compare with the baseline without increasing the training cost.
-When further increasing the number of epochs, our method continues to show certain improvements.
-Comparison using only one MLLM.
-We test the rewriting of captions using only one MLLM and utilize the generated captions for training.
-Captions are rewritten using MiniGPT-4, Otter, Qwen-VL, and LLaVA-1.5, and the results are presented in the Table 5.
-The results indicates that there is an upper limit to the performance improvement achievable by rewriting with only one model.
-Combining multiple MLLMs proves to be more effective in enhancing visual-language representation learning.
-Figure 9: Cosine similarity distribution of images and text. Our method improves the average matching degree of image-text pairs.
-Figure 10: Distribution of images and text. Our method aligns images and text better than the original distribution.
+Table 6: The retrieval performance of models trained on captions rewritten using one MLLM.
+MLLM’s Independent Effect.
+We also evaluate the performance of rewriting captions using only one MLLM and utilize the generated captions for training.
+Captions are rewritten using MiniGPT-4, Otter, Qwen-VL, and LLaVA-1.5 respectively, and the results are presented in Table 6.
+The results indicate that there is an upper limit to the performance improvement achievable by rewriting with only one MLLM.
+Besides, different MLLMs also have different strengths in knowledge.
+Combining multiple MLLMs’ knowledge proves to be more effective in enhancing visual-language pre-training.
 4.5 Visualization
-Image and text distribution visualization.
-We visualize the feature distribution of image-text pairs from the original CC3M and ours dataset in the Fig.
-10.
+Image Captioning Visualization.
+In Figure 7 we illustrate the difference in image captioning between models pre-trained on the original CC3M dataset and our improved dataset.
+Without any fine-tuning, our model exhibits a significant improvement in the ability to recognize visual concepts in images.
+(a) Image-text distribution visualization.
+(b) Cosine similarity visualization.
+Figure 8: (a) Distribution of images and texts in the original CC3M dataset and our enhanced CC3M dataset.
+Our method aligns images and texts better in the latent space.
+(b) Cosine similarity distribution of images and texts.
+Our method improves the average matching degree of image-text pairs.
+Image-text Distribution Visualization.
+We visualize the feature distribution of image-text pairs from the original CC3M and our dataset in Figure 7(a).
 It can be observed that the original CC3M dataset contains numerous discrete and unmatched image-text pairs.
@@
 This distribution perspective explains how our method aligns more continuous images with discrete language, thereby enhancing visual-language representation learning.
-Image captioning visualization.
-In the Fig.
-8 we illustrate the difference in image captioning between models pre-trained using the original CC3M dataset and our improved dataset.
-without any fine-tuning, our model exhibits a significant improvement in the ability to recognize concepts in images.
-We provide more examples of image captioning in the appendix.
 Cosine Similarity Visualization.
-We visualize the cosine similarity comparison between the original CC3M dataset and our dataset in the Fig.
-9.
-We use a pre-trained CLIP to calculate the cosine distance between images and texts.
-It is evident that our dataset exhibits a higher average similarity than the original distribution.
+We visualize the cosine similarity comparison between the original CC3M dataset and our enhanced CC3M dataset in Figure 7(b).
+We use a pre-trained CLIP to calculate the cosine distance between image embeddings and text embeddings.
+Our method exhibits a higher average similarity, which proves the effectiveness in improving the matching degree of the image-text pairs and establishing robust image-text associations for datasets.
+4.6 Comparison with More Methods
+Table 7: The image-text retrieval results on COCO and Flickr30K and classification results on ImageNet and ImageNetV2 [50] using the pre-trained CLIP on CC3M.
+Method  COCO (R@1) Flickr30k (R@1) ImageNet ImageNetV2 Avg.
+I2T T2I I2T T2I
+CLIP 13.9 9.6 26.3 18.0 14.6 12.5 15.8
+VeCLIP [29] 32.0 22.1 57.2 36.5 20.7 17.9 31.1
+CLIP 8.7 7.1 17.4 13.9 11.9 10.3 11.6
+Ours 35.9 26.5 63.5 49.3 25.0 21.4 36.9
+We also compare our work with the contemporaneous work VeCLIP [29] in the same setting in Table 7.
+We use pre-trained CLIP to compare results on image-text retrieval and image classification.
+It can be seen that our method achieves better performance improvement than the caption fusion used in VeCLIP.
+This also implies that establishing more continuous image-text correspondences is the key to improving visual-language pre-training on small-scale datasets.
+We also compare our method with  [42] in the appendix.
 5 Conclusion and Discussion
-In this paper, we propose to augment visual-language representation learning by leveraging MLLMs.
-While retaining the rich visual information of the original dataset, we utilize multiple MLLMs to extend captions for each image.
-Additionally, we introduce “text shearing” to address issues of hallucinations and intrinsic caption style in synthetic captions.
-Validated across various VLP frameworks and datasets, our method significantly improves performance on numerous downstream tasks.
-This encourages further exploration in the utilization of MLLMs in future.
-Limitations and future works Although our method has achieved remarkable results in enhancing visual-language representation learning, a certain proportion of noise still persists due to unreliable MLLM’s outputs.
-These noises limit further improvement of model’s performance.
-Future research could explore leveraging more powerful MLLMs and expanding to larger datasets.
+In this paper, we propose to augment visual-language representation learning by leveraging multiple MLLMs.
+While retaining the rich visual information of the original dataset, we utilize multiple MLLMs to extend diverse captions for each image.
+Additionally, we introduce “text shearing" to address issues of hallucinations and monotonous language style in synthetic captions.
+Validated across various visual-language pre-training frameworks and datasets, our method significantly improves performance on numerous downstream tasks.
+This encourages further exploration in the utilization of MLLMs in the future.
+Limitations and future works Although our method has achieved remarkable results in enhancing visual-language representation learning, a certain proportion of noise persists due to unreliable MLLMs’ outputs.
+These noises limit the further improvement of the model’s performance.
+Future research could explore leveraging more powerful MLLMs to generate accurate captions and expanding to larger datasets.